OpenStudy (anonymous):

The lateral area of a cube is 576 square kilometers. What is the side length?

OpenStudy (anonymous):

do u mean each lateral area or total araea

OpenStudy (anonymous):

Yes. The lateral area.

OpenStudy (anonymous):

the LSA is \[4a^{2}\] so a=12

OpenStudy (anonymous):

if its only one suface then the length is sqrt(576)=24 km but if the given figure represents total surface area then length of each side =sqrt(576/4)=sqrt(144)=12

OpenStudy (anonymous):

it is clearly mentioned lateral surface area